# Tropicalization of curve arrangement complements and arroids

### v2

@@ 4.2.2. Rank two arroids
Now we must verify balancing at each ray. There are two types of rays to verify: the $\rho_{\mathbf{p}}$ and the $\rho_{i}$. Picking a ray $\rho_{\mathbf{p}}$, we must verify that the sum $\sum_{\sigma_{i,\mathbf{p}}\succ\rho_{\mathbf{p}}}\omega(\sigma_{i,\mathbf{p}})v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ is a vector contained in $\rho_{\mathbf{p}}$, where $v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ together with the primitive vector $v_{\mathbf{p}}$ of $\rho_{\mathbf{p}}$ generates $\sigma_{i,\mathbf{p}}$. One such $v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ is precisely $[e_{j}]$, the primitive vector of $\rho_{i}$, and so we have
@@
Now by the Bézout condition, each vector $e_{j}\neq e_{i}$ appears exactly $d_{i}d_{j}$ times, so the first term is equivalent to a multiple of $[e_{i}]$ in the quotient $\mathbb{R}^{|\mathcal{A}|}/\langle(d_{1},\dots,d_{n})\rangle$. This implies in particular that the whole sum is equal to some multiple of $[e_{i}]$, and hence contained in $\rho_{i}$. ∎
Proof of Theorem 4.10.
By the constructions made in Section 4.2.1 and Section 4.2.2, we have associated a fan $\Sigma_{\mathbf{A}}$ for each transversal arroid $\mathbf{A}$, which we call the arroid fan of $\mathbf{A}$. Moreover, by the discussion in Section 4.2.1 and Lemma 4.11, these fans are tropical. ∎
@@ 4.3. Operations on transversal arroid fans
Moreover, it follows from the construction that the divisor along which the modification is performed is the reduced star ${{\Sigma_{\mathbf{A}}}^{\rho_{i}}}$, which is $\Sigma_{\mathbf{A}/i}$ by Proposition 4.13, so that we have $\Sigma_{\mathbf{A}}=\operatorname{TM}(\Sigma_{\mathbf{A}\setminus i},\Sigma_{\mathbf{A}/i})$. Moreover, the weights of $\Sigma_{\mathbf{A}}$ are then given in terms of the weights of the rays of $\Sigma_{\mathbf{A}/i}$, which is compatible with the construction of both fans as above. ∎
@@ 4.4. Tropicalization
We now show that the data gathered in the transversal arroid of an arrangement of curves is sufficient to compute its tropicalization. It is in fact sufficient to note that, all the data used in the description of the tropicalization given in Section 3 is recorded in the arroid. Our aim is instead to show that, for an arrangement where all intersections are transverse, the fan described in Section 4.2 is a fan structure on the tropicalization of the complement of the arrangement.
@@
We show that the arroid fan is supported on the tropicalization of the complement of the arrangement. Blowing up $\mathbb{P}_{K}^{2}$ in any point where more than $3$ of the curves in $\mathcal{B}$ meet, we obtain a simple normal crossing divisor $D$ which compactifies the complement $X_{\mathcal{B}}\coloneqq\mathbb{P}_{K}^{2}\setminus\cup_{C\in\mathcal{B}}\;C$. Therefore, using the map to the intrinsic torus, Theorem 2.2 gives us that the tropicalization is equal to the fan whose rays are $\operatorname{cone}([\operatorname{val}_{D}])$ in $N_{\mathbb{R}}$ for each irreducible component $D$ of the simple normal crossing divisor $\mathbf{D}$, and whose two-dimensional cones are $\operatorname{cone}([\operatorname{val}_{D}],[\operatorname{val}_{D^{\prime}}])$ for boundary divisors $D$ and $D^{\prime}$ such that $D\cap D^{\prime}\neq\emptyset$.
There are two types of irreducible components of $\mathbf{D}$: the strict transforms of the curves $C$ in $B$, whose divisorial valuations are such that the associated vectors $[\operatorname{val}_{C}]$ is a standard basis vector of $N_{\mathbb{R}}$, and the exceptional divisors $E$ of the blown-up loci $\mathbf{p}$, for which the valuation $\operatorname{val}_{E}$ computes the order of vanishing at $\mathbf{p}$. Therefore, the vector $[\operatorname{val}_{E}]$ is the sum $\sum_{i_{k}}[\operatorname{val}_{D_{i_{k}}}]$ for the $D_{i_{k}}$ which intersect at $\mathbf{p}$. In particular, it follows from this description of the rays that the tropicalization is exactly equal to the fan of the arroid $\mathbf{A}_{\mathcal{B}}$. ∎
@@ 5. Tropical homology manifold arroid fans
In this section, we study when a transversal arroid fan is a tropical homology manifold. We will first show that all tropical Borel–Moore homology of a transversal arroid fan is concentrated in top-degree, and hence torsion-free. Then we will use an Euler characteristic argument to show that being a tropical homology manifold reduces to a balancing condition. This property is a main component in understanding when the complement of an arrangement of curves is cohomologically tropical, see Theorem 6.2.
@@ 5.1. Tropical cohomology of transversal arroid fans
By Proposition 4.14, each arroid fan $\Sigma_{\mathbf{A}}$ is the tropical modification of any of its deletions, by adding a divisor to the ground set. Hence it suffices to show that the property holds for the modification of an arroid fan. Let $\Sigma_{\mathbf{A}}$ be an arroid fan, and $\operatorname{TM}(\Sigma_{\mathbf{A}},\operatorname{tropdiv}(\phi))$ a modification of $\Sigma_{\mathbf{A}}$. As pointed out in the proof of [20, Proposition 5.5], there is a short exact sequence of complexes
@@
and moreover, by Lemma 5.3, the long exact sequence in homology takes the form
@@
00{0}${{H_{p-1,0}(\operatorname{tropdiv}(\phi))}}$${{H_{p,0}(\operatorname{TM}(\Sigma,\operatorname{tropdiv}(\phi)))}}$${{H_{p,0}(\Sigma)}}$${{0,}}$00{0}${{H^{p,0}(\Sigma)}}$${{H^{p,0}(\operatorname{TM}(\Sigma,\operatorname{tropdiv}(\phi)))}}$${{H^{p-1,0}(\operatorname{tropdiv}(\phi))}}$${{0,}}$$\scriptstyle{\gamma}$$\scriptstyle{\delta}$$\scriptstyle{\delta^{\vee}}$$\scriptstyle{\gamma^{\vee}}$
@@
The [20, Lemma 5.7] holds for arbitrary tropical modifications by the above Lemma 5.2, and thus the proof given for [20, Proposition 5.6] generalizes to the present context. ∎
@@
By Lemma 5.1, the tropical Borel–Moore homology groups are such that $H_{p,q}^{BM}(\Sigma_{\mathbf{A}})=0$ for $q\neq 2$, for all $p$. Moreover, for each cone $\gamma$ of $\Sigma$, the reduced star fan ${{\Sigma}^{\gamma}}$ has $H_{p,q}^{BM}({{\Sigma}^{\gamma}})=0$ for $q\neq 1$, for all $p$, since ${{\Sigma}^{\gamma}}$ is one-dimensional. Therefore, the result follows by [1, Theorem 5.10]. ∎
@@
Since $\Sigma_{\mathbf{A}}$ is a tropical homology manifold, the reduced star along the ray corresponding to $i$, which by Proposition 4.13 is the arroid fan $\Sigma_{\mathbf{A}/i}$ of the contraction $\mathbf{A}/i$, satisfies Tropical Poincaré Duality. By Proposition 4.14, we have that $\Sigma_{\mathbf{A}}=\operatorname{TM}(\Sigma_{\mathbf{A}\setminus i},\Sigma_{\mathbf{A}/i})$, so that for each $p=0,1,2$, [20, Diagram 5.6] takes the form
00{0}${{H^{p,0}(\Sigma_{\mathbf{A}\setminus i})}}$${{H^{p,0}(\Sigma_{\mathbf{A}})}}$${{H^{p,0}(\Sigma_{\mathbf{A}/i})}}$00{0}00{0}${{H_{2-p,2}^{BM}(\Sigma_{\mathbf{A}\setminus i})}}$${{H_{2-p,2}^{BM}(\Sigma_{\mathbf{A}})}}$${{H_{2-p,2}^{BM}(\Sigma_{\mathbf{A}/i})}}$${{0.}}$$\scriptstyle{\frown[\Sigma_{\mathbf{A}\setminus i}]}$$\scriptstyle{\frown[\Sigma_{\mathbf{A}/i}]}$$\scriptstyle{\frown[\Sigma_{\mathbf{A}}]}$
By assumption both the middle and rightmost vertical arrows are isomorphisms, and therefore so is the leftmost vertical arrow. Therefore $\Sigma_{\mathbf{A}\setminus i}$ satisfies tropical Poincaré duality, and so it remains to show that each reduced star ${{\Sigma_{\mathbf{A}\setminus i}}^{\gamma}}$ of a cone $\gamma\in\Sigma_{\mathbf{A}\setminus i}$ satisfies tropical Poincaré duality. In light of Lemma 5.1, the result follows by applying [1, Proposition 5.7]. ∎
@@ 5.2. Unique balancing of arroid rays
We now begin a preliminary investigation into when the rays of an arroid fan are uniquely balanced. In light of Theorem 5.4, establishing such conditions would be equivalent to describing which arroid fans are tropical homology manifolds.
@@
holds if and only if $\alpha=\alpha_{i}$ for all $i$, where $\nu_{\rho}=\sum_{j\in\mathbf{p}}e_{j}$. Therefore the only balancing along such an edge is given by the weight $w(\mathbf{p})$ for each facet containing $\rho_{\mathbf{p}}$, as in Section 4.2. ∎
@@
holds, where the sum is taken over the set of unique points $\mathcal{P}^{u}$. Using the balancing already exhibited in Section 4.2, we may increment both sides such that $\alpha_{\mathbf{p}}\geq 0$ for all $\mathbf{p}$.
@@
Figure 4 displays an arrangement where, for the line $f$, the sets of clusters are $\{e,c,d\},\{k,h,i\}$ and the singleton $\{g\}$. The set $\{g,d\}$ is not a cluster for $f$.
@@
In Figure 5, we consider the conic $C$, which has a unique maximal cluster, comprised of all the lines together with the two additional conics $C^{\prime}$ and $C^{\prime\prime}$. Here $C^{\prime}$ is a source, $C^{\prime\prime}$ is a reservoir, and the line $L$ is an aqueduct. Therefore the unique maximal cluster is a balancing supply system.
@@
We follow the strategy of weight propagation suggested above. Let $\mathcal{C}$ be a maximal cluster, and $C^{\prime}$ a source. Since the arrangement is transverse, there are four intersection points $\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3}$ and $\mathbf{p}_{1}$ in $C\cap C^{\prime}$, and given that there are 5 lines passing through these points, they yield four distinct terms $\alpha_{\mathbf{p}}[\nu_{\mathbf{p}}]$ in the left hand sum (3). The balancing condition implies that the $\alpha_{\mathbf{p}}$ must all be equal, as otherwise the vectors $[e_{j}]$ corresponding to the lines do not appear an equal number of times. Next, let $C^{\prime\prime}$ be a reservoir, and $\mathbf{p}\in C^{\prime\prime}\cap C$ a point where there is an aqueduct $L$ between $C^{\prime}$ and $C^{\prime\prime}$. Then if the weights $\alpha_{\mathbf{p}_{i}}=\alpha$ are fixed, since the coordinate $[e_{L}]$ must appear $2\alpha$ times, we have that $\alpha_{\mathbf{p}}=\alpha$ as well. Then for the same reason, lines passing through two intersection points of $C^{\prime\prime}\cap C$ have components appearing $2\alpha$ times, thus all points of $C^{\prime\prime}\cap C$ must have the same weight. This reasoning will then also apply to any other reservoir connected to $C^{\prime\prime}$ by an aqueduct, and therefore the weights of all points in a cluster which is a balancing supply system are equal. Moreover, the balancing condition implies that the weights of two distinct clusters must agree. Therefore there is a unique balancing along the ray $\rho_{i}$. ∎
@@ 6. Cohomologically tropical arrangements
In this section, we study two properties of transverse very affine arrangements of curves. First, we show that under certain conditions, the complement of the arrangement is wunderschön in the sense of [2, Definition 1.2], which is primarily a restriction on the Mixed Hodge structure of the cohomology of this complement and of the involved curves. Next we use the study of arroid fans which are tropical homology manifold as discussed in Section 5, to show that for certain arrangements, the complement is cohomologically tropical in the sense of [2, Definition 1.1], i.e. that one may compute all the cohomology of the complement using only the tropical variety.
@@
00{0}${{\oplus_{p}\;H^{0}(p)}}$${{\oplus_{D}\;H^{2}(D)}}$${{H^{4}(\overline{X}_{\mathcal{B}})}}$${{0,}}$00{0}${{\oplus_{D}\;H^{0}(D)}}$${{H^{2}(\overline{X}_{\mathcal{B}})}}$${{0.}}$
Here the $D$’s are the components of the boundary divisors compactifying $X_{\mathcal{B}}$ to $X_{\mathcal{B}}$, and the $p$’s correspond to their pairwise intersections. Since the dual graph of the compactifying divisor is connected, the top complex only has cohomology on the left by [16, Theorem 3.1]. Moreover, the homology $H_{2}(\overline{X}_{\mathcal{B}})$ is generated by the homology of $\mathbb{P}_{\mathbb{C}}^{2}$ along with the homology of the exceptional divisors of the blown-up points [14, p. 474], so the map in the lower complex is surjective, and gives the cohomology $\mathrm{Gr}_{2}^{W}H^{2}(X_{\mathcal{B}})=0$. We therefore have $\mathrm{Gr}_{i}^{W}H^{k}(X_{\mathcal{B}})$ is 00 if $i\neq 2k$, i.e. the mixed Hodge structure is pure of weight $2k$. ∎
@@
By Proposition 6.1, this complement is necessarily wunderschön. Therefore, the equivalence described in [2, Theorem 6.1] reduces to an equivalence between $X_{\mathcal{B}}$ being cohomologically tropical and $\operatorname{trop}(X_{\mathcal{B}})=\Sigma_{\mathbf{A}_{\mathcal{B}}}$ being a tropical homology manifold. The result then follows from Theorem 5.4. ∎

### v3

@@ 4.2.2. Rank two arroids
Now we must verify balancing at each ray. There are two types of rays to verify: the $\rho_{\mathbf{p}}$ and the $\rho_{i}$. Picking a ray $\rho_{\mathbf{p}}$, we must verify that the sum $\sum_{\sigma_{i,\mathbf{p}}\succ\rho_{\mathbf{p}}}\omega(\sigma_{i,\mathbf{p}})v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ is a vector contained in $\rho_{\mathbf{p}}$, where $v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ together with the primitive vector $v_{\mathbf{p}}$ of $\rho_{\mathbf{p}}$ generates $\sigma_{i,\mathbf{p}}$. One such $v_{\sigma_{i,\mathbf{p}}/\rho_{\mathbf{p}}}$ is precisely $e_{j}]$, the primitive vector of $\rho_{i}$, and so we have
@@
Now by the Bézout condition, each vector $e_{j}\neq e_{i}$ appears exactly $d_{i}d_{j}$ times, so the first term is equivalent to a multiple of $e_{i}]$ in the quotient $\mathbb{R}^{|\mathcal{A}|}/\langle(d_{1},\dots,d_{n})\rangle$. This implies in particular that the whole sum is equal to some multiple of $e_{i}]$, and hence contained in $\rho_{i}$. ∎
Proof of Theorem˜4.10.
By the constructions made in Section˜4.2.1 and Section˜4.2.2, we have associated a fan $\Sigma_{\mathbf{A}}$ for each transversal arroid $\mathbf{A}$, which we call the arroid fan of $\mathbf{A}$. Moreover, by the discussion in Section˜4.2.1 and Lemma˜4.11, these fans are tropical. ∎
@@ 4.3. Operations on transversal arroid fans
Moreover, it follows from the construction that the divisor along which the modification is performed is the reduced star ${\Sigma_{\mathbf{A}}}^{\rho_{i}}$}, which is $\Sigma_{\mathbf{A}/i}$ by Proposition˜4.13, so that we have $\Sigma_{\mathbf{A}}=\operatorname{TM}(\Sigma_{\mathbf{A}\setminus i},\Sigma_{\mathbf{A}/i})$. Moreover, the weights of $\Sigma_{\mathbf{A}}$ are then given in terms of the weights of the rays of $\Sigma_{\mathbf{A}/i}$, which is compatible with the construction of both fans as above. ∎
@@ 4.4. Tropicalization
We now show that the data gathered in the transversal arroid of an arrangement of curves is sufficient to compute its tropicalization. It is in fact sufficient to note that, all the data used in the description of the tropicalization given in Section˜3 is recorded in the arroid. Our aim is instead to show that, for an arrangement where all intersections are transverse, the fan described in Section˜4.2 is a fan structure on the tropicalization of the complement of the arrangement.
@@
We show that the arroid fan is supported on the tropicalization of the complement of the arrangement. Blowing up $\mathbb{P}_{K}^{2}$ in any point where more than $3$ of the curves in $\mathcal{B}$ meet, we obtain a simple normal crossing divisor $D$ which compactifies the complement $X_{\mathcal{B}}\coloneqq\mathbb{P}_{K}^{2}\setminus\cup_{C\in\mathcal{B}}\;C$. Therefore, using the map to the intrinsic torus, Theorem˜2.2 gives us that the tropicalization is equal to the fan whose rays are $\operatorname{cone}([\operatorname{val}_{D}])$ in $N_{\mathbb{R}}$ for each irreducible component $D$ of the simple normal crossing divisor $\mathbf{D}$, and whose two-dimensional cones are $\operatorname{cone}([\operatorname{val}_{D}],[\operatorname{val}_{D^{\prime}}])$ for boundary divisors $D$ and $D^{\prime}$ such that $D\cap D^{\prime}\neq\emptyset$.
There are two types of irreducible components of $\mathbf{D}$: the strict transforms of the curves $C$ in $B$, whose divisorial valuations are such that the associated vectors $\operatorname{val}_{C}]$ is a standard basis vector of $N_{\mathbb{R}}$, and the exceptional divisors $E$ of the blown-up loci $\mathbf{p}$, for which the valuation $\operatorname{val}_{E}$ computes the order of vanishing at $\mathbf{p}$. Therefore, the vector $\operatorname{val}_{E}]$ is the sum $\sum_{i_{k}}[\operatorname{val}_{D_{i_{k}}}]$ for the $D_{i_{k}}$ which intersect at $\mathbf{p}$. In particular, it follows from this description of the rays that the tropicalization is exactly equal to the fan of the arroid $\mathbf{A}_{\mathcal{B}}$. ∎
@@ 5. Tropical homology manifold arroid fans
In this section, we study when a transversal arroid fan is a tropical homology manifold. We will first show that all tropical Borel–Moore homology of a transversal arroid fan is concentrated in top-degree, and hence torsion-free. Then we will use an Euler characteristic argument to show that being a tropical homology manifold reduces to a balancing condition. This property is a main component in understanding when the complement of an arrangement of curves is cohomologically tropical, see Theorem˜6.2.
@@ 5.1. Tropical cohomology of transversal arroid fans
By Proposition˜4.14, each arroid fan $\Sigma_{\mathbf{A}}$ is the tropical modification of any of its deletions, by adding a divisor to the ground set. Hence it suffices to show that the property holds for the modification of an arroid fan. Let $\Sigma_{\mathbf{A}}$ be an arroid fan, and $\operatorname{TM}(\Sigma_{\mathbf{A}},\operatorname{tropdiv}(\phi))$ a modification of $\Sigma_{\mathbf{A}}$. As pointed out in the proof of [20, Proposition 5.5], there is a short exact sequence of complexes
@@
and moreover, by Lemma˜5.3, the long exact sequence in homology takes the form
@@
[…]
@@
The [20, Lemma 5.7] holds for arbitrary tropical modifications by the above Lemma˜5.2, and thus the proof given for [20, Proposition 5.6] generalizes to the present context. ∎
@@
By Lemma˜5.1, the tropical Borel–Moore homology groups are such that $H_{p,q}^{BM}(\Sigma_{\mathbf{A}})=0$ for $q\neq 2$, for all $p$. Moreover, for each cone $\gamma$ of $\Sigma$, the reduced star fan ${\Sigma}^{\gamma}$} has $H_{p,q}^{BM}({{\Sigma}^{\gamma}})=0$ for $q\neq 1$, for all $p$, since ${\Sigma}^{\gamma}$} is one-dimensional. Therefore, the result follows by [1, Theorem 5.10]. ∎
@@
Since $\Sigma_{\mathbf{A}}$ is a tropical homology manifold, the reduced star along the ray corresponding to $i$, which by Proposition˜4.13 is the arroid fan $\Sigma_{\mathbf{A}/i}$ of the contraction $\mathbf{A}/i$, satisfies Tropical Poincaré Duality. By Proposition˜4.14, we have that $\Sigma_{\mathbf{A}}=\operatorname{TM}(\Sigma_{\mathbf{A}\setminus i},\Sigma_{\mathbf{A}/i})$, so that for each $p=0,1,2$, [20, Diagram 5.6] takes the form
[…]
By assumption both the middle and rightmost vertical arrows are isomorphisms, and therefore so is the leftmost vertical arrow. Therefore $\Sigma_{\mathbf{A}\setminus i}$ satisfies tropical Poincaré duality, and so it remains to show that each reduced star ${\Sigma_{\mathbf{A}\setminus i}}^{\gamma}$} of a cone $\gamma\in\Sigma_{\mathbf{A}\setminus i}$ satisfies tropical Poincaré duality. In light of Lemma˜5.1, the result follows by applying [1, Proposition 5.7]. ∎
@@ 5.2. Unique balancing of arroid rays
We now begin a preliminary investigation into when the rays of an arroid fan are uniquely balanced. In light of Theorem˜5.4, establishing such conditions would be equivalent to describing which arroid fans are tropical homology manifolds.
@@
holds if and only if $\alpha=\alpha_{i}$ for all $i$, where $\nu_{\rho}=\sum_{j\in\mathbf{p}}e_{j}$. Therefore the only balancing along such an edge is given by the weight $w(\mathbf{p})$ for each facet containing $\rho_{\mathbf{p}}$, as in Section˜4.2. ∎
@@
holds, where the sum is taken over the set of unique points $\mathcal{P}^{u}$. Using the balancing already exhibited in Section˜4.2, we may increment both sides such that $\alpha_{\mathbf{p}}\geq 0$ for all $\mathbf{p}$.
@@
Figure˜4 displays an arrangement where, for the line $f$, the sets of clusters are $\{e,c,d\},\{k,h,i\}$ and the singleton $\{g\}$. The set $\{g,d\}$ is not a cluster for $f$.
@@
In Figure˜5, we consider the conic $C$, which has a unique maximal cluster, comprised of all the lines together with the two additional conics $C^{\prime}$ and $C^{\prime\prime}$. Here $C^{\prime}$ is a source, $C^{\prime\prime}$ is a reservoir, and the line $L$ is an aqueduct. Therefore the unique maximal cluster is a balancing supply system.
@@
We follow the strategy of weight propagation suggested above. Let $\mathcal{C}$ be a maximal cluster, and $C^{\prime}$ a source. Since the arrangement is transverse, there are four intersection points $\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3}$ and $\mathbf{p}_{1}$ in $C\cap C^{\prime}$, and given that there are 5 lines passing through these points, they yield four distinct terms $\alpha_{\mathbf{p}}[\nu_{\mathbf{p}}]$ in the left hand sum (3). The balancing condition implies that the $\alpha_{\mathbf{p}}$ must all be equal, as otherwise the vectors $e_{j}]$ corresponding to the lines do not appear an equal number of times. Next, let $C^{\prime\prime}$ be a reservoir, and $\mathbf{p}\in C^{\prime\prime}\cap C$ a point where there is an aqueduct $L$ between $C^{\prime}$ and $C^{\prime\prime}$. Then if the weights $\alpha_{\mathbf{p}_{i}}=\alpha$ are fixed, since the coordinate $e_{L}]$ must appear $2\alpha$ times, we have that $\alpha_{\mathbf{p}}=\alpha$ as well. Then for the same reason, lines passing through two intersection points of $C^{\prime\prime}\cap C$ have components appearing $2\alpha$ times, thus all points of $C^{\prime\prime}\cap C$ must have the same weight. This reasoning will then also apply to any other reservoir connected to $C^{\prime\prime}$ by an aqueduct, and therefore the weights of all points in a cluster which is a balancing supply system are equal. Moreover, the balancing condition implies that the weights of two distinct clusters must agree. Therefore there is a unique balancing along the ray $\rho_{i}$. ∎
@@ 6. Cohomologically tropical arrangements
In this section, we study two properties of transverse very affine arrangements of curves. First, we show that under certain conditions, the complement of the arrangement is wunderschön in the sense of [2, Definition 1.2], which is primarily a restriction on the Mixed Hodge structure of the cohomology of this complement and of the involved curves. Next we use the study of arroid fans which are tropical homology manifold as discussed in Section˜5, to show that for certain arrangements, the complement is cohomologically tropical in the sense of [2, Definition 1.1], i.e. that one may compute all the cohomology of the complement using only the tropical variety.
@@
[…]
Here the $D$’s are the components of the boundary divisors compactifying $X_{\mathcal{B}}$ to $X_{\mathcal{B}}$, and the $p$’s correspond to their pairwise intersections. Since the dual graph of the compactifying divisor is connected, the top complex only has cohomology on the left by [16, Theorem 3.1]. Moreover, the homology $H_{2}(\overline{X}_{\mathcal{B}})$ is generated by the homology of $\mathbb{P}_{\mathbb{C}}^{2}$ along with the homology of the exceptional divisors of the blown-up points [14, p. 474], so the map in the lower complex is surjective, and gives the cohomology $\mathrm{Gr}_{2}^{W}H^{2}(X_{\mathcal{B}})=0$. We therefore have $\mathrm{Gr}_{i}^{W}H^{k}(X_{\mathcal{B}})$ is 0 if $i\neq 2k$, i.e. the mixed Hodge structure is pure of weight $2k$. ∎
@@
By Proposition˜6.1, this complement is necessarily wunderschön. Therefore, the equivalence described in [2, Theorem 6.1] reduces to an equivalence between $X_{\mathcal{B}}$ being cohomologically tropical and $\operatorname{trop}(X_{\mathcal{B}})=\Sigma_{\mathbf{A}_{\mathcal{B}}}$ being a tropical homology manifold. The result then follows from Theorem˜5.4. ∎